# UniChange: Unifying Change Detection with Multimodal Large Language Model

### v1

@@ Abstract
Change detection (CD) is a fundamental task for monitoring and analyzing land cover dynamics. While recent high performance models and high quality datasets have significantly advanced the field, a critical limitation persists. Current models typically acquire limited knowledge from single-type annotated data and cannot concurrently leverage diverse binary change detection (BCD) and semantic change detection (SCD) datasets. This constraint leads to poor generalization and limited versatility. The recent advancements in Multimodal Large Language Models (MLLMs) introduce new possibilities for a unified CD framework. We leverage the language priors and unification capabilities of MLLMs to develop UniChange, the first MLLM-based unified change detection model. UniChange integrates generative language abilities with specialized CD functionalities. Our model successfully unifies both BCD and SCD tasks through the introduction of three special tokens: [T1], [T2], and [CHANGE]. Furthermore, UniChange utilizes text prompts to guide the identification of change categories, eliminating the reliance on predefined classification heads. This design allows UniChange to effectively acquire knowledge from multi-source datasets, even when their class definitions conflict. Experiments on four public benchmarks (WHU-CD, S2Looking, LEVIR-CD+, and SECOND) demonstrate SOTA performance, achieving IoU scores of 90.41, 53.04, 78.87, and 57.62, respectively, surpassing all previous methods. The code is available at https://github.com/Erxucomeon/UniChange.
@@ 1 Introduction
Change Detection (CD) is the process of observing and analyzing multi-temporal remote sensing images to identify changes in land surface cover. It serves as a fundamental task in the remote sensing field and a cornerstone of modern geospatial analysis. Change detection plays a vital role in numerous applications, ranging from sustainable urban planning [41] and natural disaster assessment [51] to ecological monitoring [38] and land resource management [1]. Depending on the required level of detail, CD tasks are typically categorized into two primary sub-domains: Binary Change Detection (BCD) and Semantic Change Detection (SCD). BCD, the simpler task, primarily identifies the locations where change has occurred, but may focus on a single object of interest or not specify the nature of these transformations. While useful for automated monitoring, it provides relatively coarse-grained information. In contrast, SCD is a more complex and refined task. It not only determines the location of change but also identifies the “from-to” semantic transition [31] (e.g., from “forest” to “urban land”). This granular understanding of semantic shifts offers richer insights, providing more robust decision support for the aforementioned applications.
In the deep learning era, methods such as FCNN [9] and IFN [46] establish the Siamese Network architecture as the dominant paradigm for BCD. Subsequent research shifted focus toward enhancing the fusion and interaction of dual-temporal features. BiT [4] inserts interaction modules between the encoder and decoder to fuse dual-temporal features. Concurrently, Transformer-based architectures [2, 47, 17, 43] are proposed to extract and fuse these features at various stages within the encoder. Changer [16] pioneer a dual-temporal feature exchange paradigm by inserting a series of feature interaction layers directly within the feature extractor. Furthermore, recent works like SAM-CD [12] have begun to adapt Vision Foundation Models (VFMs) [25, 50, 35] for the BCD task. However, a fundamental limitation of these methods is the necessity of training a specialized model for each individual dataset. RSBuilding [40] partially mitigate this issue by leveraging a VFM, enabling joint training on multiple building datasets for both building extraction and change detection. However, its framework is essentially tailored to building-related tasks and cannot be extended to other categories of land cover change.
Compared to BCD, SCD is considered a more challenging task. Its core complexity lies in the fact that SCD must not only localise changes but also identify specific “from-to” semantic transitions between land cover categories. To address this challenge, the early HRSCD [10] adopt a decoupling strategy, jointly training semantic segmentation and binary change detection, which is subsequently refined by MTSCD-Net [7]. Subsequent studies have turned to deeper feature interaction and temporal modelling. Bi-SRNet [11] improves semantic change detection via bi-temporal reasoning and consistency modeling. Furthermore, SCD-SAM [31] leverages the representational capabilities of VFMs and designs a refined dual-encoder and dual-decoder architecture. Therefore, the specificity and complexity of the SCD task requires the design of a specialised architecture, which objectively hinders its harmonisation with the BCD task in terms of model architecture.
@@
By analyzing the previous BCD and SCD methods, we have identified two problems that have persisted in the field. The first is dataset incompatibility. As shown in Fig. 1, a positive sample in one dataset (e.g., ‘building’) may be explicitly defined as a negative background sample in another dataset (e.g., ‘vegetation’). This semantic conflict prevents traditional models from being jointly trained on such multi-source data. Secondly, the different requirements of BCD and SCD lead to incompatible model architectures, as shown in Fig. 2(a) and Fig. 2(b). This has led to a proliferation of highly specialized models, each with limited knowledge from a single annotation type of data (BCD or SCD). This mutually independent design approach results in poor model generalization and severely limits the versatility of the CD system in practical applications.
To address the fundamental challenges within the change detection domain, we propose UniChange. UniChange is a unified change detection model based on multimodal large language models. Not only does UniChange simultaneously support both BCD and SCD tasks, but it also enables joint training using multi-source datasets. The architecture of UniChange is shown in Fig. 2(c). We introduce three special tokens [T1], [T2], and [CHANGE] to enable the model to handle both BCD and SCD tasks concurrently. Furthermore, UniChange leverages textual cues to guide change region classification, thereby eliminating the reliance on predefined classification heads found in traditional approaches. This design facilitates the extraction of unified knowledge from diverse, multi-source change detection datasets, even when semantic category conflicts exist between datasets. Overall, our contributions are as follows:
@@
    We propose UniChange, the first unified framework for change detection based on MLLM. It is designed to simultaneously accommodate both Binary Change Detection (BCD) and Semantic Change Detection (SCD) tasks within a single end-to-end model, whilst enabling training using multi-source change detection datasets.
@@ 2.1 Change Detection
The field of Change Detection (CD) is broadly divided into two sub-fields: Binary Change Detection (BCD) and the more complex Semantic Change Detection (SCD). The latter is more challenging as it requires identifying the ”from-to” semantic information, not just the presence of change.
Early deep learning methods for BCD are dominated by CNN-based architectures. FCNN [9] is one of the early methods to utilize the paradigm of comparing features from two weight-sharing encoders. IFN [46] extend this by adding deep supervision and attention mechanisms. To better model global context, Transformer-based methods are introduced, including BiT [4], ChangeFormer [2], SwinSUNet [47] and Changer [16]. Other notable works, such as DMINet [17], and HATNet [43], further explore hybrid architectures and deep feature interaction. More recently, VFM-based methods have emerged. SAM-CD [12] focuses on adapting VFMs [25, 50, 35] to the BCD task, while RSBuilding [40] leverages VFMs to unify building-specific extraction and change detection tasks. These methods are difficult to jointly train on diverse and multi-class change detection datasets, and their generalization ability is limited.
Compared with BCD task, SCD task is more complex. HRSCD [10] and MTSCD-Net [7] utilize multi-task frameworks with separate branches for semantic segmentation and binary change detection. SCDNet [32] employs siamese networks and further utilizes attention mechanisms to refine multi-scale difference features, and Bi-SRNet [11] proposes cross-temporal semantic reasoning blocks to enhance feature consistency. Most recently, VFM-based adaptation emerges, with SCD-SAM [31] adapting VFMs by designing a specialized dual-encoder and dual-decoder system specifically for the SCD task. The differences between BCD and SCD task lead to inconsistency between their model architectures, resulting in poor versatility.
@@ 2.2 Multimodal Large Language Model
Initial MLLMs, such as LLaVA [29], MiniGPT-4 [53], and InstructBLIP [8], focus on tasks such as image captioning and visual question answering (VQA). Subsequent research has further developed this capability, enabling it to support refined visual understanding, perception and generation at both regional and pixel levels [49, 45, 33, 27, 28, 42]. While transformative, their application to remote sensing (RS) has been challenging. RS-specific MLLMs such as RSGPT [23], GeoChat [26], and GeoPixel [36] have emerged, but these largely focus on single-image interpretation. This architectural’s focus on single images makes them ill-suited for comparative, bi-temporal analysis, leaving a significant gap in their ability to perform pixel-level grounding for change detection.
Figure 3: The overview structure of UniChange is shown in (a). UniChange generates text outputs based on text instructions and dual-temporal remote sensing images. The text outputs incorporate a series of special tokens ([T1], [T2], and [CHANGE]). Concurrently, the dual-temporal remote sensing images undergo feature extraction via the Vision Backbone. These features, alongside the embeddings corresponding to the special tokens, are fed into the Token-Driven Decoder. Ultimately, UniChange generates the masks corresponding to the text instructions. The overall structure of the Token-Driven Decoder is shown in (b). It receives dual-temporal remote sensing image features and special token embeddings, generating corresponding masks.
@@ 3 Method
In this section, we first define the unified change detection task involving dual-temporal remote sensing images and textual queries, as presented in Sec. 3.1. This is followed by a detailed presentation of the UniChange architecture in Sec. 3.2, which integrates the Multimodal Large Language Model (MLMM) for instruction grounding and the Change-Perception Vision Module for pixel-level segmentation. Finally, we detail the training strategy in Sec. 3.3, including the unified loss functions designed to manage both binary and semantic constraints.
@@ 3.1 Task Definition
We define the task of unified change detection as follows. Let the input consist of a pair of dual-temporal remote sensing images, $x_{img1},x_{img2}\in R^{H\times W\times C}$, captured at times $T_{1}$ and $T_{2}$, where $H,W,$ and $C$ represent the height, width, and channel dimensions, respectively. An accompanying text input, $x_{txt}$, provides the change query instruction. This instruction can range from a general binary query (e.g., “Please segment all regions that have changed.”) to a fine-grained semantic query (e.g., “Please segment the semantic masks of the changed regions.”). The core goal is to generate a set of masks ($\hat{M}$) that accurately adhere to the semantic instructions provided by the query $x_{txt}$. This overall process can be formulated as:
@@ 3.2 Architecture
Embedding as Change. Current Change Detection (CD) models suffer from two fundamental problems: dataset incompatibility due to semantic conflicts across different sources, and architectural divergence between specialized Binary Change Detection (BCD) and Semantic Change Detection (SCD) tasks. This fragmentation prevents joint training, leading to poor generalization and limited versatility. Therefore, we propose UniChange. This framework implements our proposed ”embedding-as-change” paradigm, which integrates generative language abilities with specialized CD functionalities, thereby achieving a true unification of multi-source datasets and multi-tasks. The pipeline of our method is illustrated in Fig. 3(a).We utilize the MLLM to encode the dual-temporal remote sensing images $x_{img1}$, $x_{img2}$ and the textual query $x_{txt}$ into a unified multimodal representation. To unify both binary and semantic change detection within a single framework, we introduce three specialized tokens into the MLLM’s vocabulary: [CHANGE], [T1], and [T2]. Guided by the textual query $x_{txt}$, the MLLM then performs autoregressive generation, producing a response sequence $y_{txt}$ that strategically places these tokens based on the specific task query. This process can be formulated as:
$y_{txt}=\mathcal{F}_{MLLM}(x_{img1},x_{img2},x_{txt}),$ (2)
where $\mathcal{F}_{\text{MLLM}}$ represents the MLLM’s generation function. When the MLLM intends to generate instructions for a specific change mask (e.g., the binary change mask or a semantic mask for $T_{1}$), its output $y_{txt}$ includes the corresponding special token (e.g., [CHANGE] or [T1]). Subsequently, we extract the final-layer hidden state embedding $h_{task}$ corresponding to the position of the task-specific token ([CHANGE], [T1], or [T2]) within $y_{txt}$. This raw embedding is then projected into the feature space compatible with the downstream vision module using a dedicated $MLP(\cdot)$ layer:
$\hat{h}_{task}=MLP(h_{task}),$ (3)
these projected sparse embeddings $\hat{h}_{task}$ serve as the dynamic, instruction-guided queries.
Concurrently, the vision backbone $\mathcal{F}_{enc}$ first extracts base dense visual features from the dual-temporal remote sensing images. It then employs a feature pyramid, constructed using max-pooling and transposed convolutional layers, to generate four levels of multi-scale representations for each temporal image. This results in two multi-scale feature sets, denoted as $\{F_{1}^{i}\}$ and $\{F_{2}^{i}\}$, where $i\in\{1,2,3,4\}$. Finally, these dual-temporal visual features, along with the projected task embeddings ${\hat{h}}_{task}$, are fed into the token driven decoder $\mathcal{F}_{dec}$ to generate the final mask $\hat{M}_{task}$. The entire process can be described as:
@@
We employ a hybrid fine-tuning strategy: parameter-efficient fine-tuning (PEFT) using LoRA [22] is applied only to the MLLM’s language decoder, while all other components of the model are fully fine-tuned.
Token Driven Decoder. The token driven decoder, $\mathcal{F}_{dec}$, is specifically architected to bridge the MLLM’s high-level instructions and the pixel-level requirements of the change detection task. Its architecture is illustrated in Fig. 3(b). This decoder processes the visual features conditioned by the task-specific query embeddings derived from the MLLM. Specifically, the projected embeddings $\{\hat{h}_{t1},\hat{h}_{t2},\hat{h}_{change}\}$ (obtained from Eq. 3) serve as the initial task queries. Let $E^{0}$ represent the concatenation of these initial queries:
@@
the concatenated query embeddings $E^{0}$ are then hierarchically refined through four decoders. Each decoder $i$ processes its corresponding multi-scale visual features $F_{1}^{i}$ and $F_{2}^{i}$ by first adding a learnable positional encoding ($PE$), flattening ($\Phi_{Flat}$) them into token sequences, and concatenating ($\Phi_{Cat}$) them into a unified visual sequence $T^{i}$:
$T^{i}=\Phi_{Cat}(\Phi_{Flat}(F_{1}^{i}+PE),\Phi_{Flat}(F_{2}^{i}+PE)).$ (6)
A dual cross-attention mechanism then refines both the task queries and the visual sequence. The task queries $E^{i-1}$ (from the previous level) first attend to themselves ($\mathcal{A}_{self}$), then attend to the visual sequence $T^{i}$ ($\mathcal{A}_{cross}$), and are subsequently processed by an FFN ($\mathcal{F}_{FFN}$). Concurrently, the visual sequence $\hat{T}^{i}$ is also updated by attending to these newly refined task queries. This recursive process is defined as:
$\begin{gathered}  E^{i}=\mathcal{F}_{FFN}(\mathcal{A}_{cross}(\mathcal{A}_{self}(E^{i-1}),T^{i})),\\   \hat{T}^{i}=\mathcal{A}_{cross}(T^{i},E^{i}),\end{gathered}$ (7)
the refined queries $E^{4}$ from the final level, along with the refined visual sequences $\hat{T}^{i}$ from all levels, are used for the final mask generation.
First, we process the refined visual sequences $\hat{T}^{i}$ (where $i\in\{1,2,3,4\}$) from each decoder level. As shown in Eq. 8, the sequence $\hat{T}^{i}$ from each level is split and reshaped ($\Phi_{Split/Reshape}$) back into its corresponding 2D dual-temporal feature maps, $\hat{F}_{1}^{i}$ and $\hat{F}_{2}^{i}$:
@@
These multi-level features are then individually upsampled ($\Phi_{Up}$) to a uniform scale and concatenated ($\Phi_{Cat}$) along the channel dimension. A fusion module ($\Phi_{fuse}$) integrates them to create two final dense representations, $F_{t1}$ and $F_{t2}$. The dense change feature map, $F_{change}$, is then derived by element-wise subtraction:
$\begin{gathered}F_{t1}=\Phi_{fuse}(\Phi_{Cat}(\{\Phi_{Up}(\hat{F}_{1}^{i})\})),\\ F_{t2}=\Phi_{fuse}(\Phi_{Cat}(\{\Phi_{Up}(\hat{F}_{2}^{i})\})),\\ F_{change}=F_{t1}-F_{t2}.\end{gathered}$ (9)
The refined query embeddings $E^{4}$ from the final decoder level are then processed. These embeddings are first projected ($\Phi_{proj}$) and then split ($\Phi_{Split}$) into the individual task representations $\hat{e}_{t1}$, $\hat{e}_{t2}$, and $\hat{e}_{change}$. These projected embeddings are used to filter their corresponding dense feature maps ($F_{t1}$, $F_{t2}$, $F_{change}$) to generate the final masks:
$\hat{e}_{t1},\hat{e}_{t2},\hat{e}_{change}$ $=\Phi_{Split}(\Phi_{proj}(E^{4})),$ (10)
$\hat{M}_{task}$ $=\mathcal{M}_{gen}(F_{task},\hat{e}_{task}),$
@@
where $\mathcal{M}_{\text{gen}}$ represents the mask generation function, which performs Einstein summation. This unified structure allows UniChange to flexibly generate the corresponding masks based on the user’s instructions.

### v2

@@ Abstract
Change detection (CD) is a fundamental task for monitoring and analysing land cover dynamics. While recent high performance models and high quality datasets have significantly advanced the field, a critical limitation persists. Current models typically acquire limited knowledge from single-type annotated data and cannot concurrently leverage diverse binary change detection (BCD) and semantic change detection (SCD) datasets. This constraint leads to poor generalisation and limited versatility. The recent advancements in Multimodal Large Language Models (MLLMs) introduce new possibilities for a unified CD framework. We leverage the language priors and unification capabilities of MLLMs to develop UniChange, the first MLLM-based unified change detection model. UniChange integrates generative language abilities with specialised CD functionalities. We introduce three special tokens: [T1], [T2], and [CHANGE], utilising their embeddings as the key to query variations. This approach successfully accommodates both BCD and SCD tasks. Furthermore, UniChange utilises text prompts to guide the identification of change categories, eliminating the reliance on predefined classification heads. This design allows UniChange to effectively acquire knowledge from multi-source datasets, even when their class definitions conflict. Experiments on four public benchmarks (WHU-CD, S2Looking, LEVIR-CD+, and SECOND) demonstrate SOTA performance, achieving IoU scores of 90.41, 53.04, 78.87, and 57.62, respectively, surpassing all previous methods. The code is available at https://github.com/Erxucomeon/UniChange.
@@ 1 Introduction
Change detection (CD) is the process of observing and analysing multi-temporal remote sensing images to identify changes in land surface cover. It serves as a fundamental task in the remote sensing field and a cornerstone of modern geospatial analysis. Change detection plays a vital role in numerous applications, ranging from sustainable urban planning [wellmann2020remote] and natural disaster assessment [zheng2021building] to ecological monitoring [singh1986change] and land resource management [asadzadeh2022uav]. Depending on the required level of detail, CD tasks are typically categorised into two primary sub-domains: binary change detection (BCD) and semantic change detection (SCD). BCD, the simpler task, primarily identifies the locations where change has occurred, but may focus on a single object of interest or not specify the nature of these transformations. While useful for automated monitoring, it provides relatively coarse-grained information. In contrast, SCD is a more complex and refined task. It not only determines the location of change but also identifies the “from-to” semantic transition [mei2024scd] (e.g., from “forest” to “urban land”). This granular understanding of semantic shifts offers richer insights, providing more robust decision support for the aforementioned applications.
@@
In the deep learning era, methods such as FC-Siam-diff [daudt2018fully] and IFN [zhang2020deeply] establish the siamese network architecture as the dominant paradigm for BCD. Subsequent research shifts focus toward enhancing the fusion and interaction of dual-temporal features. BiT [chen2022remote] fuses dual-temporal features using token-space context modeling and feature refinement. Concurrently, Transformer-based architectures [bandara2022transformer, zhang2022swinsunet, xu2024hybrid] are proposed to fuse and model long-range dependencies between these features at various stages within the encoder. Changer [fang2023changer] pioneers a dual-temporal feature exchange paradigm by inserting a series of feature interaction layers directly within the feature extractor. Furthermore, recent works like SAM-CD [ding2024adapting] have begun to adapt Vision Foundation Models (VFMs) [kirillov2023segment, zhao2023fast, ravi2024sam] for the BCD task. However, a fundamental limitation of these methods is the necessity of training a specialised model for each individual dataset. RSBuilding [wang2024rsbuilding] partially mitigates this issue by leveraging a VFM, enabling joint training on multiple building datasets for both building extraction and change detection. However, its framework is essentially tailored to building-related tasks and cannot be extended to other categories of land cover change.
Compared to BCD, SCD is considered a more challenging task. Its core complexity lies in the fact that SCD must not only localise changes but also identify specific “from-to” semantic transitions between land cover categories. To address this challenge, the early HRSCD [daudt2019multitask] adopts a decoupling strategy, jointly training semantic segmentation and binary change detection, which is subsequently refined by MTSCD-Net [cui2023mtscd]. Subsequent studies have turned to deeper feature interaction and temporal modelling. Bi-SRNet [ding2022bi] improves semantic change detection via dual-temporal reasoning and consistency modelling. Furthermore, SCD-SAM [mei2024scd] leverages the representational capabilities of VFMs and designs a refined dual-encoder and dual-decoder architecture. The SCD task’s inherent specificity and complexity necessitate a specialised architecture, thereby hindering architectural unification between BCD and SCD tasks.
@@
By analysing the previous BCD and SCD methods, we have identified two problems that have persisted in the field. The first is dataset incompatibility. As shown in Fig. 1, a positive sample in one dataset (e.g., “building”) may be explicitly defined as a negative background sample in another dataset (e.g., “vegetation”). This semantic conflict prevents traditional models from being jointly trained on such multi-source datasets. Secondly, the different requirements of BCD and SCD lead to incompatible model architectures, as shown in Fig. 2(a) and Fig. 2(b). This has led to a proliferation of highly specialised models, each with limited knowledge from a single annotation type of dataset (BCD or SCD). This mutually independent design approach results in poor model generalisation and severely limits the versatility of the CD system in practical applications.
To address the fundamental challenges within the change detection domain, we propose UniChange. UniChange is a unified change detection model based on the multimodal large language model. Not only does UniChange simultaneously support both BCD and SCD tasks, but it also enables joint training using multi-source datasets. The architecture of UniChange is shown in Fig. 2(c). We introduce three special tokens [T1], [T2], and [CHANGE] to enable the model to handle both BCD and SCD tasks concurrently. Furthermore, UniChange leverages textual cues to guide change region classification, thereby eliminating the reliance on predefined classification heads found in traditional approaches. This design facilitates the extraction of unified knowledge from diverse, multi-source change detection datasets, even when semantic category conflicts exist between datasets. Overall, our contributions are as follows:
@@
    We propose UniChange, the first unified framework for change detection based on the MLLM. It is designed to simultaneously accommodate both binary change detection (BCD) and semantic change detection (SCD) tasks within a single end-to-end model, whilst enabling training using multi-source change detection datasets.
@@ 2.1 Change Detection
The field of change detection (CD) is broadly divided into two sub-fields: binary change detection (BCD) and the more complex semantic change detection (SCD). The latter is more challenging as it requires identifying the “from-to” semantic information, not just the presence of change.
Early deep learning methods for BCD are dominated by CNN-based architectures. FC-Siam-diff [daudt2018fully] is one of the early methods to utilise the paradigm of comparing features from two weight-sharing encoders. IFN [zhang2020deeply] extends this by adding deep supervision and attention mechanisms. To better model global context, Transformer-based methods are introduced, including BiT [chen2022remote], ChangeFormer [bandara2022transformer], SwinSUNet [zhang2022swinsunet] and Changer [fang2023changer]. Other notable works, such as DMINet [feng2023change], and HATNet [xu2024hybrid], further explore hybrid architectures and deep feature interaction. More recently, VFM-based methods have emerged. SAM-CD [ding2024adapting] focuses on adapting VFMs [kirillov2023segment, zhao2023fast, ravi2024sam] to the BCD task, while RSBuilding [wang2024rsbuilding] leverages VFMs to unify building-specific extraction and change detection tasks. However, these methods are difficult to jointly train on diverse and multi-class change detection datasets, limiting their generalisation.
Compared with BCD task, SCD task is more complex. HRSCD [daudt2019multitask] and MTSCD-Net [cui2023mtscd] decouple the task into binary change detection and semantic segmentation subtasks, integrating them within a multi-task framework for refined feature association. SCDNet [peng2021scdnet] employs siamese networks and further utilises attention mechanisms to refine multi-scale difference features, while Bi-SRNet [ding2022bi] proposes cross-temporal semantic reasoning blocks to enhance feature consistency. Most recently, VFM-based adaptation emerges. SCD-SAM [mei2024scd] adapts VFMs by designing a specialised dual-encoder and dual-decoder system specifically for the SCD task. The differences between BCD and SCD tasks lead to inconsistency between their model architectures, resulting in poor versatility.
@@ 2.2 Multimodal Large Language Model
Initial MLLMs, such as LLaVA [liu2023visual], MiniGPT-4 [zhu2023minigpt], and InstructBLIP [dai2023instructblIP], focus on tasks such as image captioning and visual question answering (VQA). Subsequent research has further developed this capability, enabling it to support refined visual understanding, perception and generation at both regional and pixel levels [zhang2024gpt4roi, you2023ferret, peng2024grounding, lai2024lisa, li2025visual, wu2025representation]. While transformative, their application to remote sensing (RS) has been challenging. RS-specific MLLMs such as RSGPT [hu2025rsgpt], GeoChat [kuckreja2024geochat], and GeoPixel [shabbir2025geopixel] have emerged, but these methods largely focus on single-image interpretation. This architecture, which focuses on single images, makes it ill-suited for comparative, dual-temporal analysis. It leaves a significant gap in the ability to perform pixel-level grounding for change detection.
Figure 3: The overview structure of UniChange is shown in (a). UniChange generates text outputs based on text instructions and dual-temporal remote sensing images. The text outputs incorporate a series of special tokens ([T1], [T2], and [CHANGE]). Concurrently, the dual-temporal remote sensing images undergo feature extraction via the Vision Backbone. These features, alongside the embeddings corresponding to the special tokens, are fed into the Token Driven Decoder. Ultimately, UniChange generates the masks corresponding to the text instructions. The overall structure of the Token Driven Decoder is shown in (b). It receives dual-temporal remote sensing image features and special token embeddings, generating corresponding masks.
@@ 3 Method
In this section, we first define the unified change detection task. As detailed in Sec. 3.1, this task involves dual-temporal remote sensing images and textual queries. Subsequently, we present the UniChange architecture in detail in Sec. 3.2. This architecture comprises an MLLM for understanding changes, a Vision Backbone for extracting features from dual-temporal remote sensing images, and a specialised Token Driven Decoder. Finally, we detail the training strategy in Sec. 3.3, including the unified loss functions designed to manage both binary and semantic constraints.
@@ 3.1 Task Definition
We define the task of unified change detection as follows. Let the input consist of a pair of dual-temporal remote sensing images, $x_{img1},x_{img2}\in R^{H\times W\times C}$, captured at times $T_{1}$ and $T_{2}$, where $H,W,$ and $C$ represent the height, width, and channel dimensions, respectively. An accompanying text input, $x_{txt}$, provides the change query instruction. This instruction may be a binary change query (e.g., “Please segment all areas that have undergone change.”) or a more complex semantic change query (e.g., “Please segment the semantic masks of the changed areas.”). The core goal is to generate a set of masks ($\hat{M}$) that accurately adhere to the semantic instructions provided by the query $x_{txt}$. This overall process can be formulated as
@@ 3.2 Architecture
Embedding as Change. Current change detection (CD) models suffer from two fundamental problems: dataset incompatibility due to semantic conflicts across different sources, and architectural divergence between specialised binary change detection (BCD) and semantic change detection (SCD) tasks. This fragmentation prevents joint training, leading to poor generalisation and limited versatility. Therefore, we propose UniChange. This framework implements our proposed “embedding as change” paradigm. This paradigm integrates generative language abilities with specialised CD functionalities, thereby achieving a true unification of multi-source datasets and multi-tasks. The framework of our method is illustrated in Fig. 3(a). MLLM takes the dual-temporal remote sensing images $x_{img1}$, $x_{img2}$ and a given textual instruction $x_{txt}$ as input. To unify both binary and semantic change detection within a single framework, we introduce three specialised tokens into the MLLM’s vocabulary: [T1], [T2] and [CHANGE]. Conditioned on the textual instruction $x_{txt}$, the MLLM autoregressively generates the response sequence $y_{txt}$. These tokens are strategically placed within this sequence in accordance with the specific task query. This process can be formulated as
$y_{txt}=\mathcal{F}_{MLLM}(x_{img1},x_{img2},x_{txt}).$ (2)
When the MLLM intends to generate masks for a specific change (e.g., the binary change mask or a semantic mask for $T_{1}$ and $T_{2}$), its output $y_{txt}$ includes the corresponding special token (e.g., [T1], [T2] or [CHANGE]). Subsequently, we extract the MLLM last-layer embedding $h_{task}$ corresponding to the position of a specific token ([T1], [T2] or [CHANGE]) within $y_{txt}$. This raw embedding is then projected by a dedicated $MLP(\cdot)$ layer to make it compatible with the downstream vision module’s feature space:
$\hat{h}_{task}=MLP(h_{task}).$ (3)
These projected sparse embeddings (where each is individually generated as $\hat{h}_{task}$ via Eq. 3) serve as the dynamic, instruction-guided queries.
Concurrently, the vision backbone $\mathcal{F}_{enc}$ first extracts base dense visual features from the dual-temporal remote sensing images. A feature pyramid is then constructed using max-pooling and transposed convolutional layers. It generates four levels of multi-scale image features for each temporal image. This results in two multi-scale image feature sets, denoted as $\{F_{1}^{i}\}$ and $\{F_{2}^{i}\}$, where $i\in\{1,2,3,4\}$. Finally, these dual-temporal visual features, along with the projected task embeddings ${\hat{h}}_{task}$, are fed into the token driven decoder $\mathcal{F}_{dec}$ to generate the final mask $\hat{M}_{task}$. The entire process can be described as
@@
We employ a hybrid fine-tuning strategy. Only the language decoder of the MLLM undergoes LoRA fine-tuning [hu2022lora]. All other components of the model are fully parameterised for fine-tuning.
Token Driven Decoder. Inspired by RSBuilding, the token driven decoder, $\mathcal{F}_{dec}$, is specifically architected to translate the MLLM’s high-level instructions into pixel-level segmentation masks for the change detection task. Its architecture is illustrated in Fig. 3(b). This decoder is employed to process the dual-temporal remote sensing image features and the projected special token embeddings. Specifically, the projected embeddings $\{\hat{h}_{t1},\hat{h}_{t2},\hat{h}_{change}\}$ (obtained from Eq. 3) serve as the initial task queries. Let $E^{0}$ represent the concatenation ($\Phi_{Cat}$) of these initial queries:
@@
the concatenated task query $E^{0}$ is then hierarchically refined through four decoder layers. Each decoder layer processes its corresponding multi-scale image features $F_{1}^{i}$ and $F_{2}^{i}$ by flattening ($\Phi_{Flat}$) them into token sequence, and concatenating ($\Phi_{Cat}$) them into a unified visual sequence $T^{i}$:
$T^{i}=\Phi_{Cat}(\Phi_{Flat}(F_{1}^{i}),\Phi_{Flat}(F_{2}^{i})).$ (6)
The SAM Decoder Layer then refines both the task query and the visual sequence. The task query $E^{i-1}$ (from the previous level) first attends to itself ($\mathcal{A}_{self}$), then attends to the visual sequence $T^{i}$ ($\mathcal{A}_{cross}$), and is subsequently processed by an FFN ($\mathcal{F}_{FFN}$). Next, the visual sequence $T^{i}$ is also updated by attending to this newly refined task query ($\mathcal{A}_{cross}$). The positional encodings are added to queries and keys in all attention operations. This process is defined as
$\begin{gathered}  E^{i}=\mathcal{F}_{FFN}(\mathcal{A}_{cross}(\mathcal{A}_{self}(E^{i-1}),T^{i})),\\   \hat{T}^{i}=\mathcal{A}_{cross}(T^{i},E^{i}).\end{gathered}$ (7)
The refined task query $E^{4}$ is taken from the final level. The query and the refined visual sequences $\{\hat{T}^{i}\}$ from all levels are then used for the final mask generation.
The visual sequences $\{\hat{T}^{i}\}$ from each level are split and reshaped ($\Phi_{Split/Reshape}$) back into their corresponding 2D dual-temporal feature maps, $\{\hat{F}_{1}^{i}\}$ and $\{\hat{F}_{2}^{i}\}$:
@@
Next, three feature sets are prepared: $\{\hat{F}_{1}^{i}\}$, $\{\hat{F}_{2}^{i}\}$, and their element-wise difference $\{\hat{F}_{1}^{i}-\hat{F}_{2}^{i}\}$. All three sets are then uniformly processed by a three-step procedure: upsampling ($\Phi_{Up}$), concatenation ($\Phi_{Cat}$), and fusion ($\Phi_{fuse}$). This derives $F_{t1}$, $F_{t2}$, and $F_{change}$, respectively:
$\begin{gathered}F_{t1}=\Phi_{fuse}(\Phi_{Cat}(\{\Phi_{Up}(\hat{F}_{1}^{i})\})),\\ F_{t2}=\Phi_{fuse}(\Phi_{Cat}(\{\Phi_{Up}(\hat{F}_{2}^{i})\})),\\ F_{change}=\Phi_{fuse}(\Phi_{Cat}(\{\Phi_{Up}(\hat{F}_{1}^{i}-\hat{F}_{2}^{i})\})).\end{gathered}$ (9)
Concurrently, the refined task query $E^{4}$ from the final decoder level is split and projected ($\Phi_{Split/Proj}$) into the individual task embeddings $\hat{e}_{t1}$, $\hat{e}_{t2}$, and $\hat{e}_{change}$:
$\hat{e}_{t1},\hat{e}_{t2},\hat{e}_{change}$ $=\Phi_{Split/Proj}(E^{4}).$ (10)
Finally, these projected embeddings are used to filter their corresponding features ($F_{t1}$, $F_{t2}$, $F_{change}$) to generate the final masks:
$\hat{M}_{task}$ $=\mathcal{M}_{gen}(F_{task},\hat{e}_{task}),$ (11)
@@
where $\mathcal{M}_{\text{gen}}$ represents the mask generation function. This function performs Einstein Summation.
This unified structure allows UniChange to flexibly generate the corresponding masks based on user instructions.